Find the slope intercept form for the line satisfying the conditions.
Perpendicular to -1/4 x+2,passing through the point (7,-3)

Answer by LinnW(1048)

A line perpendicular to a line with slope -1/4 is 4
We want an equation of the form y = Mx + b .
For our problem, M = 4
Since our line needs to go through (7,-3), x = 7 and y = -3
Substituting into y = Mx + b
-3 = 4(7) + b
-3 = 28 + b
Add -28 to each side
-31 = b
So our equation is y = 4x -31